OpenStudy (anonymous):

I WILL MEDAL!! Fifty is what percent of 40? A. 20% B. 80% C. 110% D. 125%

OpenStudy (anonymous):

d

OpenStudy (anonymous):

When we say "Y is X percent of Z", we mean (X/100)*Z=Y In this case, Z is 40 and Y is 50, so: (X/100)*40=50 Dividing both sides by 40 gives: (x/100)=(50/40) Multiply both sides by 100 gives x = 100 * (50/40) = 5000/40 = 125 So 50 is 125% of 40

OpenStudy (anonymous):

Is it C?

OpenStudy (anonymous):

concluding all of that the answer shall be D.125%

OpenStudy (anonymous):

No, 110% of 40 is 44

OpenStudy (anonymous):

We can set up an equation: Fifty is x% of forty 50 = x% of 40 50 = 40x What's 50 / 40?

OpenStudy (anonymous):

@I"like"bribro Please do not give answers, it's against OpenStudy's Code-of-Conduct. We are here to guide the user to the answer, so we need his participation. http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

125%

OpenStudy (anonymous):

@KendrickLamar2014 That goes for you too..

OpenStudy (anonymous):

i explained my answer tho

OpenStudy (anonymous):

Yes, but we are still not supposed to give answers..we need the asker's participation..